Zeke is driving from his house to the beach and wants to model how far he is from home. He is currently 50 miles from the beach and will drive 65 miles per hour en route to the beach. What equation will he need to use?
Answer by Boreal(15235)

distance=time*velocity
65 miles per hour is 1.08333 miles per minute
His distance from home is d=1.05833 mi/min* t(min)

so time=d/v=50/65 units mi/mi/hour or hours
=12/13 hours or 55.4 minutes